### v3

# Rate of epidemic spreading on complex networks

Abstract
The initial phase of an epidemic is often characterized by an exponential increase in the number of infected individuals. In this Letter, we predict the exponential spreading rate of an epidemic on a complex network. We first find an expression of the reproduction number for a network, based on the degree distribution, the network assortativity, and the level of clustering. We then connect this reproduction number and the disease infectiousness to the spreading rate. Our result holds for a broad range of networks, apart from networks with very broad degree distribution, where no clear exponential regime is present. Our theory bridges the gap between classic epidemiology and the theory of complex network, with broad implications for model inference and policy making.
Epidemic modeling permits rationalization of evidences and provides crucial information for policy making. Epidemiological models should incorporate the key features that determine epidemic spreading. For example, social structures play a crucial role in epidemic spreading Buckee et al. (2021), and in the effectiveness of intervention strategies Patwardhan et al. (2023); Granell and Mucha (2018). In these social structures, the number of contacts per individual is highly variable. For instance, individuals with most contacts contributed disproportionately to the spreading of COVID-19 Kumar et al. (2020); Sneppen et al. (2020).
Social structures can be conveniently modeled by means of complex networks Newman (2018), in which nodes represent individuals and links denote pairs of individuals that are in contact. A success of network theory has been to determine the epidemic threshold on networks Pastor-Satorras and Vespignani (2001); Newman (2002); Pastor-Satorras et al. (2015), i.e., the transition point from a localized outbreak to a widespread epidemic. This theory has shown that scale-free networks are extremely vulnerable to epidemics Pastor-Satorras and Vespignani (2001); Barthélemy et al. (2005); Pastor-Satorras et al. (2015); Pósfai and Barabási (2016) . Clustering, i.e., the tendency of nodes to form tightly connected communities, also affects epidemic spreading. Many real-world networks possess both heavy-tailed degree distribution and high level of clustering Ravasz and Barabási (2003).
Another key feature of an epidemic is its infectiousness, i.e., the rate at which an infected individual infects its contacts. The infectiousness incorporates multiple factors, such as the viral load, social behavior, and possibly environmental factors Grassly and Fraser (2008). Because of changes in these factors, the infectiousness of an individual changes over time in a way that is particular to each disease. This time-dependence profoundly impacts how a disease spreads.
Social structures and individual infectiousness are crucial in determining the initial stage of an epidemic, in which most individuals are susceptible and the number of infected individuals usually grows exponentially. This stage is of central importance for novel epidemics, or during emergence of novel variants, as it often requires rapid policy making. At later stages, other factors become relevant, such as individual recovery time, their behavioral response Weitz et al. (2020), and the possible occurrence of reinfections.
Despite the vast body of theory on epidemics in networks Pastor-Satorras et al. (2015), general results on the speed of epidemic spreading are surprisingly sparse. The rate of epidemic spreading can be calculated for unclustered, uncorrelated networks, and with constant infectiousness Barthélemy et al. (2004). A recent theory estimates the average time it takes for an epidemic to reach an individual at a certain distance from patient zero Moore and Rogers (2020). The spread of an epidemic on a network can be exactly solved Merbis and Lodato (2022), but this calculation is practically feasible only for very small networks. Other approaches avoid a detailed description of social structures, and effectively model the population heterogeneity by assigning different risk propensities to individuals Rose et al. (2021); Berestycki et al. (2023).
In this Letter, we theoretically predict the exponential rate of epidemic spreading on a complex network. Our theory quantifies the impact of main properties of the network, such as assortativity and clustering, on the reproduction number of the epidemic. We further link the reproduction number and the individual infectiousness with the exponential spreading rate. Our approach leads to precise predictions for a broad range of networks.
Epidemic spreading in well-mixed populations. We begin by summarizing how the exponential spreading rate of an epidemic is determined in well-mixed populations Grassly and Fraser (2008). In this case, the cumulative number $I(t)$ of infected individuals at time $t$ is governed by the renewal equation
[…]
where the basic reproduction number $R_{0}$ represents the average number of secondary infections by each infected individual, and the dot denotes a time derivative. The function $w(\tau)$ represents the probability density that a given secondary infection occurs after a time $\tau$ from the primary one. Equation (1) is linear because of the assumption that the epidemic is in the exponential stage and thus far from saturation. We now substitute the ansatz $I(t)\sim e^{\Lambda t}$ into Eq. (1) and take the limit of large $t$, obtaining the Euler-Lotka equation
$\frac{1}{R_{0}}=\langle e^{-\Lambda\tau}\rangle\,,$ (2)
[…]
Over the course of the epidemic, each individual can be infected at most once and never formally recovers, although their infectiousness may wane over time. Our model can thus be see as a non-Markovian Susceptible-Infected model. The time dependence of $\lambda(\tau)$ is important to model real-world epidemics. For example, the infectiousness of COVID-19 markedly increases for 3–6 days and then decays for about 11 days Hakki et al. (2022), and is therefore poorly described by a Markovian model with constant $\lambda(\tau)$.
In the model, the number of infected individuals grows in time until it saturates (Fig. 1c). An alternative to the temporal perspective is to represent the epidemic in generations (Fig. 1d). We call $Z_{n}$, $n=0,1,\dots$ the average number of infected individuals in the $n$th generation. The initial generation, $n=0$, is constituted by the patient zero, $Z_{0}=1$. In this representation, the infected individuals form a tree (Fig. 1e).
Figure 1: (a) Epidemic spreading on a network. Individuals (nodes) can be infected (red) or healthy (blue). Infected individuals spread the infection at a rate $\lambda(\tau)$ to each of their neighbors, where $\tau$ is the time since they were infected. If neighbors have already been infected, transmission does not result in an infection. (b) Time-dependent spreading rate $\lambda(\tau)$ and resulting distribution $\psi(\tau)$ of infection times. (c) An epidemic trajectory. The exponential phase ends in a saturation phase, where a sizeable fraction of the individuals are infected. (d) Epidemic tree. The epidemic initiates with one infected individual (patient zero). Subsequent infections are represented by continuous lines. Dashed lines represents generations.
Reproduction number on a network. The basic reproduction number $R_{0}$ for a network is the average number of links of an infected individual in the early stage of the epidemic, excluding the one from whom the infection came from. The effective reproduction number $R$ is the average number of infections caused by an infected individual. In contrast with $R_{0}$, $R$ does not include contacts that are infected by other individuals. This difference is crucial in clustered networks.
To estimate $R$ for a network, we characterize individuals by their degree $k$. We introduce the reproduction matrix
$M_{kk^{\prime}}=(k^{\prime}-1-m_{kk^{\prime}})P(k|k^{\prime})\,,$ (3)
whose elements are the average numbers of susceptible individuals of degree $k$ connected to a node of degree $k^{\prime}$. In Eq. (3), $P(k|k^{\prime})$ is the probability that a neighbor of a node of degree $k^{\prime}$ has degree $k$ and $m_{kk^{\prime}}$ is the average number of triangles a $kk^{\prime}$-link is part of. The term proportional to $m_{kk^{\prime}}$ discounts for contacts of a given individual that were already infected by someone else, due to the presence of triangles. In Eq. (3), we neglect higher order loops such as squares. The reproduction matrix has been previously used to study percolation on networks Serrano and Boguná (2006).
We express the average number $Z_{n}$ of infected individuals at generation $n\geq 1$ in terms of $M_{kk^{\prime}}$ by summing over the degrees of individuals at each generation:
$Z_{n}=\sum_{k_{1}\dots k_{n}}M_{k_{n}k_{n-1}}M_{k_{n-1}k_{n-2}}\dots M_{k_{2}k_{1}}k_{1}P(k_{1})\,.$ (4)
At the first generation, we obtain $Z_{1}=\bar{k}$. For large $n$, we have $Z_{n}\sim R^{n}$, where we identify the reproduction number $R$ as the leading eigenvalue of $M_{kk^{\prime}}$. The associated right eigenvector $v_{k^{\prime}}$ represents the degree distribution of infected individuals sup .
In the simple case of an unclustered network $(m_{kk^{\prime}}=0$ for all links) in which the degrees of connected nodes are uncorrelated, we have
$R=R_{0}=\frac{\overline{k^{2}}-\overline{k}}{\overline{k}}\,,$ (5)
where the bar represents an average over the degree distribution. Equation (5) is a classic result Pastor-Satorras and Vespignani (2001); Callaway et al. (2000); Newman (2002). We seek to extend it to the general case of clustered and correlated networks. We quantify degree correlations between connected nodes by the assortativity coefficient $r=1/\sigma_{r}^{2}(\sum_{kk’}kk^{\prime}P(k,k’)-(\overline{k^{2}}/\overline{k})^{2})$,with $\sigma_{r}^{2}=\overline{k^{3}}/\overline{k}-(\overline{k^{2}}/\overline{k})^{2}$ sup . We find that, for small clustering and assortativity coefficients, the effective reproduction number is expressed by
$R_{\mathrm{pert}}=(1-r)(R_{0}-m_{1})+r\left(\frac{\frac{\overline{k(k-1)^{2}}}{\overline{k}}-2m_{2}+m_{1}^{2}}{R_{0}-m_{1}}\right)\,,$ (6)
where $c_{k}$ is the probability that two neighbors of a node of degree $k$ are themselves connected, $m_{1}=\overline{k(k-1)c_{k}}/\overline{k}$ is the average number of triangles a link is part of, and $m_{2}=\overline{k(k-1)^{2}c_{k}}/\overline{k}$. We obtained Eq. (6) by a perturbative expansion in $r$ sup .
We compare $R_{\mathrm{pert}}$ given by Eq. (6) with the maximum eigenvalue $R$ of the matrix $M_{kk^{\prime}}$ for different network models, see Table 1. The relative difference between $R$ and $R_{\mathrm{pert}}$ is within $2\%$ for all the models we considered, except for Barabási-Albert with parameter $m=1$ ($20.6\%$) and $m=2$ ($7.1\%$), where $m$ represents the number of nodes a new node attaches to. Approximating $R$ with $R_{0}$ as given by Eq. (5) leads to substantially larger errors, supporting that clustering and assortativity play an important role. Here and in the following, when computing $R_{\mathrm{pert}}$ using Eq. (6) and $R_{0}$ using Eq. (5) for a given network, we always interpret the averages as empirical averages. With this choice, moments of the degree distribution remain finite even in very heterogeneous networks. In practical cases, Eq. (6) predicts that $R_{\mathrm{pert}}$ increases with the assortativity $r$. Accordingly, assortativity lowers the epidemic threshold, as previously known for the unclustered case Goltsev et al. (2012). However, this dependence is usually weak, unless the degree distribution has heavy tails. In contrast, clustering substantially reduces $R_{\mathrm{pert}}$. For this reason, we find that $R_{0}$ overestimates $R$ in all the synthetic networks we considered.
Table 1: Reproduction numbers. Here, $R$ is the numerically computed leading eigenvalue of Eq. (3). The expressions of $R_{0}$ and $R_{\mathrm{pert}}$ are given by Eq. (5) and Eq. (6), respectively. To assess the impact of assortativity and clustering, we also compute $R_{\mathrm{pert}}$ for $r=0$ and for $\bar{c}=\sum_{k}c_{k}P(k)=0$. BAm: Barabási-Albert with parameter $m$. $\text{ER}_{d}$: Erdős-Rényi with mean degree $d$. WSk,p: Watts-Strogatz starting from a 1D lattice with $k$ nearest neighbors and rewiring probability $p$.
Network $R_{0}$ $R$ $R_{\mathrm{pert}}$ $R_{\mathrm{pert}}^{(r=0)}$ $R_{\mathrm{pert}}^{(\overline{c}=0)}$ $r$ $\overline{c}$
$\text{BA}_{1}$ 13.9 8.5 6.8 13.9 6.6 -0.011 0
$\text{BA}_{2}$ 20.4 15.5 14.4 20.3 14.2 -0.008 0.0001
$\text{BA}_{12}$ 85.4 80.8 79.5 83.6 80.4 -0.004 0.0003
$\text{ER}_{6}$ 6.0 6.0 6.0 6.0 6.0 0.001 0.000
$\text{WS}_{8,2}$ 7.2 4.9 4.9 4.9 7.0 -0.022 0.334
$\text{WS}_{6,2}$ 5.2 3.6 3.6 3.6 5.0 -0.029 0.315
Network Euler-Lotka equation. We now use our estimate of $R$ to predict how epidemics spread in time. The average total number ${I}_{k}(t)$ of infected individual of degree $k$ at time $t$ is governed by the renewal equation
$\dot{I}_{k}(t)=\sum_{k^{\prime}}\int_{0}^{t}\dot{I}_{k^{\prime}}(t-\tau)M_{kk^{\prime}}\psi(\tau)\mathrm{d}\tau\,.$ (7)
Equation (7) extends Eq. (1) to complex networks. Assuming exponential growth, $\dot{I}_{k^{\prime}}(u)\propto v_{k^{\prime}}e^{\Lambda t}$, we obtain for large time
$\frac{1}{R}=\left\langle e^{-\Lambda\tau}\right\rangle\,,$ (8)
[…]
We shall use Eq. (8) as an implicit relation for the unknown $\Lambda$, and call $\Lambda_{\mathrm{EL}}$ the solution of the equation. We can also explicitly express $\Lambda$ from Eq. (8), but only in an approximate way. In particular, using the first two moments of $\psi(\tau)$, we obtain
$\Lambda\approx\frac{\ln R}{\langle\tau\rangle}+\left(\ln R\right)^{2}\frac{\sigma^{2}}{2\langle\tau\rangle^{3}}\,,$ (9)
where $\sigma^{2}=\langle\tau^{2}\rangle-\langle\tau\rangle^{2}$, see sup . An alternative is to combine Eq. (8) with the Jensen inequality, that states that $\langle e^{-\Lambda\tau}\rangle\geq e^{-\Lambda\langle\tau\rangle}$. This leads to the lower bound $\Lambda\geq(\ln R)/\langle\tau\rangle$. This bound is saturated when $\psi(\tau)$ is a Dirac delta function.
Synthetic networks. The network Euler-Lotka equation successfully predicts the rate of epidemic spreading on synthetic complex networks, Fig. 2a and Fig. 2b. Our battery of tests includes a range of models characterized by different degree distribution, assortativity, and clustering. For each model, we use three different generation time distributions, see Fig. 2c. This choice qualitatively affect the epidemics. In particular, for peaked distributions, the exponential growth appears modulated by oscillations (Fig. 2a), since early generations of the epidemic are nearly synchronised. To extrapolate the leading exponential behavior from these curves, we fit them to a generalised logistic function
$I(t)=\frac{N}{\left(1+\left((N/I_{0})^{\nu}-1\right)e^{-\Lambda_{0}\nu t}\right)^{1/\nu}}\,,$ (10)
where $N$ is the size of the network and $I_{0}$, $\Lambda_{0}$, and $\nu$ are free parameters, see sup . Expanding Eq. (10) for small $t$ we obtain $\log I(t)\approx\log I_{0}+t\Lambda_{0}[1-(I_{0}/N)^{\nu}]$. This leads us to define the finite-size growth rate $\Lambda_{\rm FIT}=\Lambda_{0}(1-\eta)$ and the finite-size parameter $\eta=\left(I_{0}/N\right)^{\nu}\in[0,1]$. The parameter $\eta$ is also a measure of deviations from an exponential behavior, see sup .
Figure 2: Network Euler-Lotka equation predicts the epidemic spreading rate on synthetic networks. (a) and (b): Average trajectory $I(t)$ over 500 realizations (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed) and $\Lambda_{\text{EL}}$ (black dashed; based on $R_{\text{pert}}$, Eq. (6)) for Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) network with Gamma-distributed infection times and Watts-Strogatz network with Weibull-distributed infection times. (c) Solution of Eq. (8) compared with simulations for various networks of sizes up to $10^{6}$, see Ref. sup for details.
When computing $\Lambda_{\mathrm{EL}}$, we set $R=R_{\text{pert}}$ as defined in Eq. (6) as the reproduction number. As expected, the predictions are considerably worse if we instead use $R_{0}$, see sup . The difference between $\Lambda_{\mathrm{FIT}}$ and $\Lambda_{\mathrm{EL}}$ is within $6\%$ for all networks, see Fig. 2c. The only exception is the Barabási-Albert model with parameter $m=1$ (average error $15\%$). One explanation for this discrepancy is that the exponential regime for this model is not well defined for $N\rightarrow\infty$, since $\bar{k^{2}}$ diverges in this limit. Another explanation is that, for $m=1$, $R_{\mathrm{pert}}$ does not accurately match the leading eigenvalue of the reproduction matrix (see Table 1).
Figure 3: Network Euler-Lotka equation predicts the epidemic spreading rate on real-world networks. We obtained $I(t)$ by averaging $500$ simulations using log-normal infection times for each network from Refs. Clauset et al. (2016); Leskovec and Krevl (2014); Kunegis (2013) with at least $1000$ nodes sup . (a) and (b): Average trajectory $I(t)$ (red), $\Lambda_{\text{FIT}}$ (blue dot-dashed) and $\Lambda_{\text{EL}}$ (black dashed) for powergrid and PIN networks. (c) Exponential spreading rate $\Lambda_{\text{EL}}$ (dashed line) versus $\Lambda_{\text{FIT}}$ (markers). (d) Degree distribution tail exponent ($\gamma)$ and finite-size parameter ($\eta$) affect the estimation error $\varepsilon=2|\Lambda_{\text{FIT}}-\Lambda_{\rm EL}|/(\Lambda_{\text{FIT}}+\Lambda_{\rm EL})$. Each dot represents a network, dashed lines mark the scale-free region $2<\gamma\leq 3$. (e) and (f): Average degree $\bar{k}_{n}$ of infected individuals at $n$-th generation for (a) Erdős-Rényi ($N=10^{6}$, $\overline{k}=4$) and (b) Youtube (scale-free, $\gamma\approx 2.2$). Dashed lines represent power-law fits for $n\in[1,6]$.
Real-world networks. We apply the network Euler-Lotka equation to a large set of social, biological, technological, and transportation networks. These real-world networks often exhibit more complex structures than synthetic networks, with tightly linked communities, strong degree-correlation and broad tails in the degree distribution. Although they constitute a challenging test, the network Euler-Lotka equation holds well, see Fig. 3a and Fig. 3b for examples.
More generally, we find that $\Lambda_{\text{EL}}$ well matches $\Lambda_{\text{FIT}}$ whenever the finite-size parameter $\eta$ detects a clear exponential regime, see Fig. 3c. Although some of these networks are highly clustered, using $R_{\mathrm{pert}}$ as expressed by Eq. (6) does not perform worse than $R$ sup . In Fig. 3c, we have omitted three road networks, for which the fit of Eq. (10) fails to converge. However, a function of the form $\dot{I}(t)\sim t^{\beta}$ well fits the epidemic trajectories in these cases sup , likely because these networks are embedded in a two-dimensional physical space (see Moore et al. (2024)).
For networks in which $\eta<0.5$, the predictions of Eq. (8) using $R_{\text{pert}}$ are more accurate than those using $R_{0}$. However, for larger values of $\eta$, $R_{0}$ performs slightly better on average sup . Even in these cases, $R_{\mathrm{pert}}$ provides a better approximation of $R$ sup . The advantage of $R_{0}$ for large $\eta$ might be due to a compensatory effect: using $R_{0}$ tends to overestimate $R$, while the network Euler-Lotka equation tends to underestimate $\Lambda$ compared to $\Lambda_{\mathrm{FIT}}$.
All of the networks where the prediction error is large are scale-free and characterized by a large value of $\eta$, see Fig. 3d. Here, scale-free means that the tail of the degree distribution scales as $k^{-\gamma}$, with $\gamma\leq 3$, see Voitalov et al. (2019) and sup . In contrast, the error is weakly affected by clustering sup . The error for scale-free networks can be explained by the early infection of hubs, which depletes the tail of the degree distribution. In networks that are not scale-free, we find that the average degree of infected individual is constant among generations as expected (Fig. 3e). In contrast, we find this average degree to decay as a power law in scale-free networks (Fig. 3f), leading to a non-steady spreading.
Discussion. In this Letter, we have related the exponential spreading rate of an epidemic on a complex network with the infectiousness and the network structure. Our theory accurately predicts the spreading rate on a wide range of synthetic and real-world networks, aside from scale-free networks where the exponential regime is hindered by finite-size effects.
When analyzing the early stage of real epidemics, departure from an exponential spreading is often interpreted as a consequence of behavioral response or containment measures (see, e.g., Maier and Brockmann (2020)). Our results show that non-exponential epidemic trajectories are common in heavy-tailed networks, thus providing an alternative explanation for these observations.
The classic Euler-Lotka equation is often used to estimate $R$ in real epidemics. However, this approach is often inaccurate, because of difficulties in estimating both $\Lambda$ and the infectiousness distribution (see, e.g., Lauer et al. (2020); Ferretti et al. (2020); Nishiura et al. (2020)). For this reason, approaches based on contact tracing have been developed Kojaku et al. (2021); Birello et al. (2024). Our theory paves the way to use basic structural information on the interaction networks to directly estimate $R$.
We have focused for simplicity on static networks. Epidemic dynamics are more complex in temporal networks Cai et al. (2024), for example due bursty social interactions Tkachenko et al. (2021). Interactions could be also weighted, corresponding to contacts of different duration and intensity Ferretti et al. (2024). Our approach is flexible enough to be extended to these cases.
Acknowledgements.
We thank C. Castellano and I. Neri for comments on a preliminary version of this manuscript. We thank Joshua Weitz for discussions. We are grateful for the help and support provided by the Scientific Computing and Data Analysis section of Core Facilities at OIST.